A refrigerator with a coefficient of performance of 1.80 absorbs 3.15×104 J of heat from the low-temperature reservoir during each cycle.

How much mechanical work is required to operate the refrigerator for a cycle?

How much heat does the refrigerator discard to the high-temperature reservoir during each cycle?
Answer #1

Solutiowo Given! Coefficient of peeformance, k = 1.80 1 Absorbs heat from the cold reservoir Qc = 3.15*10ts ① let wo be the m>> Q4 = (3.15 x 10t + x 104 + 1.75X106) J = |QH = 4.9x10ts ... Heat discard to hot reservoir during each cycle iů 4.9x107]Please give positive response it helps me alot....thank you